EQUAL volumes of  0.1N AgNO3 and 0.2M NaCl are mixed.What will be the concentration of NO3- ions in the mixture?

Grade:9

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

The concentration of NO3 will be 0.2 M

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sree ram
13 years ago

    [NO3]- in 0.1M AgNO3=0.1M

   When mixed with equal volume of 0.2M NaCl Solution,

   The volume of the solution is double.So Concentration of NO3- is halved.

    [NO3-]=0.1/2=0.05M

Raghunath Yadav Karike
13 years ago

Dear ajesh,

In this the NO3- ions are coming from the compound AgNO3.And its normality aswell as molarity will be 0.1 in magnitude.

When it free its(NO3-) concentration = 0.1 M

But the equal amount of another is added , we know that concentration is inversely proportional to volume. When volume is doubled means concentration will be halved.So required concentration = 0.1/2=0.05

All the best.